Grade 12th passInorganic Chemistry

Sir I don`t know how to do these type of questions plz help me out to do by showing me the steps:-What is effective atomic no. of Mn in [Mn(CO)5]- ?

sohan goswami
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
We can write it as Mn2(CO)10
 

The formula for effective atomic number(EAN) is => Z - O.S + 2CN

here,

Z=atomic number

O.S=oxidation state

CN=coordination number

So, FOR 2 Mn

Z=2×25

O.S=0

CN=10 (for 10 CO ) + 1 (for 1 Mn-Mn sigma bond)

Thus CN = 11

EAN = 2×25 - 0 + 11×2

50+ 22

72

Thus 72 for 2 Mn

or 72/2 for 1Mn

= 36